John was 10 years older than Jane 10 years ago. Jim was 5 years younger than John 10 years ago. If Jane is 25 now what is the present age of John and Jim?
Answer by mananth(16946)

John was 10 years older than Jane 10 years ago. Jim was 5 years younger than John 10 years ago. If Jane is 25 now what is the present age of John and Jim?
...
Jane -------25 years now
..
10 years ago jane - 15 years
..
John 10 years ago was 10 years older than Jane ----15+10=25 years
...
Jim ------ 5 years younger than John 10 years ago -------20 years
..
Jim = 10+20=30 years now
John = 25+10 = 35 years now.